# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 4.2.2 RL for Iterative Code Refinement
A second line of research targets more complex coding tasks that require debugging and iterative refinement. In these scenarios, an agent may need multiple attempts to produce correct code, using feedback (e.g., error messages or failed test results) to incrementally improve its solution, which is closer to real-world tasks.
@@
[…]
PSGPO [40] Process - -
[…]
PRLCoder [280] Process CodeT5+, Unixcoder, T5-base -
[…]
Focused-DPO [285] Process DeepSeek-Coder-6.7B-Base/Instruct, Magicoder-S-DS-6.7B, Qwen2.5-Coder-7B-Instruct -
[…]
Qwen Team [277] Outcome Qwen-2.5-Coder-Instruct-7B/32B -
RL for Iterative Code Refinement
RLEF [286] Outcome Llama-3.0-8B-Instruct, Llama-3.1-8B/70B-Instruct -
[…]
IterPref [289] Process Deepseek-Coder-7B-Instruct, Qwen2.5-Coder-7B, StarCoder2-15B -
LeDex [290] Process StarCoder-15B, CodeLlama-7B/13B -
[…]
ReVeal [292] Process DAPO-Qwen-32B -
Posterior-GRPO [282] Process Qwen2.5-Coder-3B/7B-Base, Qwen2.5-Math-7B -
[…]
@@
[…]
Golubev et al. [299] Process Qwen2.5-72B-Instruct -
[…]
@@
Process-supervised approaches explicitly guide how the model debugs. IterPref [289] constructs localized preference pairs from iterative debugging traces and applies targeted preference optimization to penalize faulty regions, improving correction accuracy with minimal collateral updates. LeDex [290] couples explanation-driven diagnosis with self-repair: it automatically curates explanation–refinement trajectories and applies dense, continuous rewards to jointly optimize explanation quality and code correctness via PPO, yielding consistent pass@1 gains over SFT-only coders. Beyond explanation-driven shaping, some works like CTRL [291] explicitly train separate critic models to evaluate each attempted refinement and provide gradient signals to the policy, though at the cost of added inference overhead. Process-gated schemes such as Posterior-GRPO [282] and reliability-enhancing filters [283] further stabilize learning by ensuring that intermediate reasoning helps only when the final output is validated. ReVeal [292] is a self-evolving code agent that iteratively creates tests to verify its own code, employing per-turn rewards to deepen reasoning and enhance recovery from errors.
@@ 4.2.3 RL for Automated Software Engineering
The frontier of LLM-based coding agents involves full-scale software engineering tasks that are long-horizon and require tool use, planning, and reasoning across multiple steps. These include tasks like reading and modifying a large codebase, implementing a new feature or fixing a complex bug via multiple edits, using external tools (compilers, linters, version control, shells) and verifying the results over many iterations. Such scenarios demand an agentic approach. RL fine-tuning has proven uniquely effective at instilling long-term decision-making skills that static supervised training cannot capture [293]. Recent open-source research has begun to tackle this level of complexity with RL-based training frameworks.
@@
End-to-end training in realistic environments demonstrates that sparse—but validated—success signals can scale. DeepSWE performs large-scale RL on software engineering missions using verified task completion as the sole reward, achieving leading open-source results on SWE-bench–style evaluations [293]. SWE-RL extracts rule-based, outcome-oriented signals from GitHub commit histories, enabling training on authentic improvement patterns and generalization to unseen bug-fixing tasks [294]. Satori-SWE introduces an evolutionary RL-enabled test-time scaling method (EvoScale) that trains models to self-improve generations across iterations for sample-efficient software engineering tasks [295]. OS-R1 [298] presents a rule-based reinforcement learning framework for Linux kernel tuning, enabling efficient exploration, accurate configuration, and superior performance over heuristic methods. RLCoder frames retrieval-augmented repository-level code completion as an RL problem, using perplexity-based feedback to train a retriever to fetch helpful context without labeled data [296]. Qwen3-Coder performs large-scale execution-driven reinforcement learning on long-horizon, multi-turn interactions across 20,000 parallel environments, yielding state-of-the-art performance on benchmarks like SWE-Bench Verified [181]. In machine learning domains, ML-Agent executes multi-step pipelines (e.g., automated ML), optimizing performance-based terminal rewards [297].
@@
Dense supervision over agentic trajectories improves credit assignment across many steps. From the optimization perspective, long-context, multi-turn software agents benefit from stabilized policy-gradient variants; e.g., Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) improves training stability and performance on SWE-bench Verified through multi-turn code generation and debugging interactions, leveraging long-context feedback [299]. SWEET-RL trains multi-turn agents on ColBench (backend and frontend tasks), leveraging privileged information during RL to reduce exploration noise and improve long-horizon generalization [145].
@@
Commercial systems such as OpenAI’s Codex and Anthropic’s Claude Code have emphasized preference-aligned fine-tuning and reinforcement learning to improve usefulness and safety in code generation and editing workflows [300, 301]. While concrete training details are limited publicly, these systems underscore the growing role of RL in aligning agentic behavior with developer-centric objectives in practical IDE and terminal environments.
@@ 4.3 Mathematical Agent
We note that RLVR methods such as DAPO [169], GRPO [302], and GRESO [303] have consistently played a substantial role in recent enhancements of mathematical reasoning in LLMs. However, given their broader relevance across reasoning tasks, we discuss them in Section 2.7, instead of elaborating here.
@@ 4.3.1 RL for Informal Mathematical Reasoning
Informal mathematics essentially refers to reasoning and expression in natural language. Such reasoning may incorporate symbols, such as variables (e.g., $x$, $A_{i}$), equations (e.g., $x+y=z$), or function names, but no finite and explicit set of logical rules defines their syntactic validity, and no formal semantics precisely determines their interpretation and meaning [304, 305].
While informal mathematical reasoning relaxes strict rigor at the detail level, it affords greater expressive flexibility and better captures the high-level structure of arguments. This makes it particularly suited for a variety of math tasks such as mathematical word problem solving, equation manipulation, and symbolic computation [91, 306]. Although general-purpose programming languages are symbolic, they lack the rigor and formal semantics of proof-assistant languages, and are therefore regarded as informal when applied to mathematical reasoning [304], typically through tool invocation of executors such as Python with numerical or symbolic libraries.
@@
Outcome-only methods define rewards solely by final numerical or symbolic correctness (e.g. algebraic equations) during RL. Empirically, such training often leads to emergent agentic behaviors, including adaptive tool use interleaved with natural language reasoning. ARTIST [91] introduces a framework for tool-integrated agentic reasoning, interleaving tool invocations, e.g. code execution, directly within the reasoning chain. Trained with outcome-only rewards, it achieves strong performance and observes emergent agentic behaviors, including self-reflection, and context-aware CoT, which further shows that by integrating dynamic tool use with RL, agentic tool-integrated reasoning could learn optimal strategies for interacting with environments, highlighting the potential of RL to internalize tool-integrated reasoning strategies in LLMs. Similarly, ToRL [92] improves performance by exploiting the scaling of tool-integrated reasoning RL and encouraging code execution behaviour, and experiments show emergent cognitive behaviors, such as adaptive tool-use, self-correction based on tool feedback, and adaptive computational reasoning. ZeroTIR [306] investigates the scaling law of RL from outcome-based rewards for Tool-Integrated Reasoning with Python code execution settings, revealing a strong correlation between training computational effort and the spontaneous code execution frequency, the average response length, and the final task accuracy, which corroborates the empirical emergence of tool-integrated reasoning strategies. TTRL [151] leverages majority voting to estimate rewards, enabling training on unlabeled data. Fine-tuned on these majority-vote rewards, it not only surpasses the base model’s maj@n accuracy but also achieves an empirical performance curve and upper bound that, surprisingly, closely approach those of direct RL training with labeled test answers on MATH-500, underscoring its practical value and potential. However, RENT [307] suggests that majority voting is short on generalization, it applies only to questions with deterministic answers, and will not work on free-response outputs. To address this limitation, it extends the entropy minimization idea [308] to RL, using the token-level average negative entropy as a reward to guide learning, achieving improvements on an extensive suite of benchmarks including math problem solving, suggesting that confidence-based reward shaping can serve as a path toward continual improvement. Alternatively, Satori [295] proposes Chain-of-Action-Thought (COAT), a variant of CoT that explicitly integrates action choices, and modularizes reasoning into 3-fold meta-actions, including continuation, reflection, and exploration of alternatives, and internalizes this behavior via RL with outcome-only rewards. In particular, 1-shot RLVR [309] studies data efficiency of outcome-only RL with verifier signals. Surprisingly, they found that RL with only 1 example performs close to using a 1.2k-example dataset, and with 2 examples comes close to using the 7.5k MATH training dataset. They also highlight an intriguing phenomenon, named post-saturation generalization, that test accuracy continues to improve even after the training accuracy on the single example approaches 100%. In addition to correctness, hallucination remains a major challenge in informal mathematical reasoning, motivating methods that explicitly promote trustworthiness. For instance, Kirchner et al. [310] propose a game-theoretic training algorithm that jointly optimizes for both correctness and legibility. Inspired by Prover-Verifier Games [311], the method alternates between training a small verifier that predicts solution correctness, a "helpful" prover that generates solutions accepted by the verifier, and a "sneaky" prover that aims to fool it. Empirically, this increases the helpful prover accuracy, verifier robustness and legibility (measured by human accuracy in time-constraint verification tasks). This result suggests that verifier-guided legibility optimization can enhance the interpretability and trustworthiness of LLM-generated informal reasoning. Recent rStar2-Agent [107] is a 14B-parameter math reasoning model trained with agentic reinforcement learning using a high-throughput Python execution environment, a novel GRPO-RoC algorithm to resample on correct rollouts amid tool-noise, and a multi-stage training recipe—achieving state-of-the-art results in just 510 RL steps, achieving average pass@1 scores of 80.6% on AIME24 and 69.8% on AIME25.
@@
Process-aware methods leverage intermediate evaluators (e.g. unit tests, assertions, sub-task checks) to provide denser feedback, shaping credit assignment and improving tool-integrated reasoning (TIR). START [312] guides TIR by injecting handcrafted hint text into Long CoT traces, typically after conjunction words or before the CoT stop token, to encourage code executor calls during inference. This enables test-time scaling that improves reasoning accuracy. The collected trajectories are then used to fine-tune the model, internalizing the tool-invocation behavior. LADDER [313] introduces a training-time framework where an LLM recursively generates and solves progressively simpler variants of a complex problem, using verifiable reward signals to guide a difficulty-based curriculum, and achieves substantial improvements in mathematical reasoning. An additional test-time RL step (TTRL) further enhances performance. The authors suggest that this approach of self-generated curriculum learning with verifiable feedback may generalize beyond informal mathematical tasks to any domain with reliable automatic verification. To improve performance on complex problems, SWiRL [314] synthesizes step-wise tool use reasoning data by iteratively decomposing solutions, and then adopts a preference-based step-wise RL approach to fine-tune the base model on the multi-step trajectories. While many of these approaches exploits inference-time interventions, they often suffer from generalization limitations due to their reliance on manually designed logical structures. To overcome this, RLoT [315] instead trains a lightweight navigator agent model with RL to adaptively enhance reasoning, showing improved generalization across diverse tasks.
While informal approaches excel at word problems and symbolic computations, they struggle to extend effectively to advanced mathematical tasks such as automated theorem proving. This limitation arises from two fundamental challenges: evaluation difficulty, which demands machine-verifiable feedback unavailable to informal methods, and scarcity of high-quality formal proof data. [304, 305]

### v4

@@ 4.2.2 RL for Iterative Code Refinement
A second line of research targets more complex coding tasks that require debugging and iterative refinement. In these scenarios, an agent may need multiple attempts to improve solutions, using feedback from human requirements or failed test results, which is closer to real-world tasks.
@@
Figure 7: Benchmark Performance of RL-Enhanced Code and SWE Methods. Scores are pass@1 unless otherwise specified.
@@
AceCoder (Zeng et al., 2025b) Outcome Qwen2.5-Coder-7B-Base/Instruct, Qwen2.5-7B-Instruct GitHub
DeepCoder-14B (Luo et al., 2025c) Outcome DeepSeek-R1-Distilled-Qwen-14B GitHub
RLTF (Liu et al., 2023b) Outcome CodeGen-NL 2.7B, CodeT5-770M GitHub
CURE (Wang et al., 2025q) Outcome Qwen2.5-7B/14B-Instruct, Qwen3-4B GitHub
Absolute Zero (Zhao et al., 2025a) Outcome Qwen2.5-7B/14B, Qwen2.5-Coder-3B/7B/14B, Llama-3.1-8B GitHub
StepCoder (Dou et al., 2024) Process DeepSeek-Coder-Instruct-6.7B GitHub
PSGPO (Dai et al., 2025) Process Qwen2.5-Coder-7B-Instruct -
CodeBoost (Wang et al., 2025n) Process Qwen2.5-Coder-7B-Instruct, Llama-3.1-8B-Instruct, Seed-Coder-8B-Instruct, Yi-Coder-9B-Chat GitHub
PRLCoder (Ye et al., 2025b) Process CodeT5+, Unixcoder, T5-base -
o1-Coder (Zhang et al., 2024d) Process DeepSeek-1.3B-Instruct GitHub
Posterior-GRPO (Fan et al., 2025a) Process Qwen2.5-Coder-3B-Base, Qwen2.5-Coder-7B-Instruct, Qwen2.5-Math-7B -
Policy Filtration for RLHF (Zhang et al., 2025c) Process DeepSeek-Coder-6.7B, Qwen1.5-7B GitHub
CodeFavor (Liu et al., 2024a) Process Mistral-NeMo-12B-Instruct, Gemma-2-9B-Instruct, Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3 GitHub
Focused-DPO (Zhang et al., 2025o) Process DeepSeek-Coder-6.7B-Base/Instruct, Magicoder-S-DS-6.7B, Qwen2.5-Coder-7B-Instruct -
Re:Form (Yan et al., 2025a) Outcome Qwen2.5 (0.5B–14B) GitHub
Qwen Team (Feng et al., 2025c) Outcome Qwen2.5-Coder-7B/32B-Instruct -
RL for Iterative Code Refinement
RLEF (Gehring et al., 2025) Outcome Llama-3.0-8B-Instruct, Llama-3.1-8B/70B-Instruct -
@@
R1-Code-Interpreter (Chen et al., 2025h) Outcome Qwen2.5-7B/14B-Instruct-1M, Qwen2.5-3B-Instruct GitHub
IterPref (Wu et al., 2025c) Process Deepseek-Coder-7B-Instruct, Qwen2.5-Coder-7B, CodeQwen1.5-7B-Chat, StarCoder2-15B -
LeDex (Jiang et al., 2024) Process StarCoder-15B, CodeLlama-7B/13B -
CTRL (Xie et al., 2025) Process Qwen2.5-Coder-7B/14B/32B-Instruct GitHub
ReVeal (Jin et al., 2025c) Process DAPO-Qwen-32B -
@@
DeepSWE (Luo et al., 2025b) Outcome Qwen3-32B GitHub
SWE-RL (Wei et al., 2025e) Outcome Llama-3.3-70B-Instruct GitHub
Satori-SWE (Zeng et al., 2025a) Outcome Qwen2.5-Coder-32B-Instruct GitHub
RLCoder (Wang et al., 2025p) Outcome CodeLlama-7B, StarCoder-7B, StarCoder2-7B, DeepSeekCoder-1B/7B GitHub
Qwen3-Coder (Team, 2025b) Outcome Qwen3-Coder-480B-A35B-Instruct GitHub
ML-Agent (Liu et al., 2025q) Outcome Qwen2.5-7B-Base/Instruct, DeepSeek-R1-Distill-Qwen-7B GitHub
OS-R1 (Lin et al., 2025b) Outcome Qwen2.5-3B/7B-Instruct GitHub
Golubev et al. (2025) Process Qwen2.5-72B-Instruct -
SWEET-RL (Zhou et al., 2025e) Process Llama-3.1-8B/70B-Instruct GitHub
@@
Process-supervised approaches explicitly guide how the model debugs. IterPref (Wu et al., 2025c) constructs localized preference pairs from iterative debugging traces and applies targeted preference optimization to penalize faulty regions, improving correction accuracy with minimal collateral updates. LeDex (Jiang et al., 2024) couples explanation-driven diagnosis with self-repair: it automatically curates explanation–refinement trajectories and applies dense, continuous rewards to jointly optimize explanation quality and code correctness via PPO, yielding consistent pass@1 gains over SFT-only coders. Beyond explanation-driven shaping, some works like CTRL (Xie et al., 2025) explicitly train separate critic models to evaluate each attempted refinement and provide gradient signals to the policy, though at the cost of added inference overhead. ReVeal (Jin et al., 2025c) extends process-level refinement into a self-evolving agent that autonomously generates tests and learns from per-turn rewards to enhance reasoning and recovery from errors.
@@ 4.2.3 RL for Automated Software Engineering
End-to-end training in realistic environments demonstrates that sparse—but validated—success signals can scale. DeepSWE performs large-scale RL on software engineering missions using verified task completion as the sole reward, achieving leading open-source results on SWE-bench–style evaluations (Luo et al., 2025b). SWE-RL extracts rule-based, outcome-oriented signals from GitHub commit histories, enabling training on authentic improvement patterns and generalization to unseen bug-fixing tasks (Wei et al., 2025e). Satori‑SWE introduces an evolutionary RL-enabled test-time scaling method (EvoScale) that trains models to self-improve generations across iterations for sample-efficient software engineering tasks (Zeng et al., 2025a). OS-R1 (Lin et al., 2025b) presents a rule-based reinforcement learning framework for Linux kernel tuning, enabling efficient exploration, accurate configuration, and superior performance over heuristic methods. RLCoder frames retrieval-augmented repository-level code completion as an RL problem, using perplexity-based feedback to train a retriever to fetch helpful context without labeled data (Wang et al., 2025p). Qwen3‑Coder performs large-scale execution-driven reinforcement learning on long-horizon, multi-turn interactions across 20,000 parallel environments, yielding state-of-the-art performance on benchmarks like SWE‑Bench Verified (Team, 2025b). In machine learning domains, ML-Agent executes multi-step pipelines (e.g., automated ML), optimizing performance-based terminal rewards (Liu et al., 2025q).
@@
Dense supervision over agentic trajectories improves credit assignment across many steps. From the optimization perspective, long-context, multi-turn software agents benefit from stabilized policy-gradient variants; e.g., Decoupled Clip and Dynamic sAmpling Policy Optimization (DAPO) improves training stability and performance on SWE-bench Verified through multi-turn code generation and debugging interactions, leveraging long-context feedback (Golubev et al., 2025). SWEET-RL trains multi-turn agents on ColBench (backend and frontend tasks), leveraging privileged information during RL to reduce exploration noise and improve long-horizon generalization (Zhou et al., 2025e).
@@
Commercial systems such as OpenAI’s Codex and Anthropic’s Claude Code have emphasized preference-aligned fine-tuning and reinforcement learning to improve usefulness and safety in code generation and editing workflows (OpenAI, 2025a; Anthropic, 2025). While concrete training details are limited publicly, these systems underscore the growing role of RL in aligning agentic behavior with developer-centric objectives in practical IDE and terminal environments.
4.2.4 Emerging Paradigms
Code World Models
A recent paradigm shift departs from traditional neural approximations by framing the world model itself as executable code. In these Code World Models (CWMs), agents synthesize programs to explicitly define transition and reward dynamics, enabling model-based planning via verifiable, symbolic simulation rather than opaque latent states.
GIF-MCTS (Dainese et al., 2024) formulates world-model construction as program induction: the LLM edits an “Environment” class and a search procedure selects versions that best explain offline transitions, yielding executable models suitable for downstream planning. WorldCoder (Tang et al., 2024) represents dynamics and rewards as explicit Python functions and refines them through an iterative synthesize–repair process guided by transition consistency and optimism constraints. Meta’s 32B CWM (team et al., 2025) strengthens this paradigm by providing an open-weights model trained on interpreter traces and agentic trajectories to improve program synthesis and execution fidelity. Recent work further applies CWMs to general game environments (Lehrach et al., 2025), where an LLM induces complete rule-based simulators and planning is performed directly on the executable model.
Implementing these programmatic world-model paradigms often incurs substantial inference cost, since agents repeatedly synthesize, refine, and query executable simulators. In practice, low-bit quantization (e.g., 8-bit for workstation GPUs like RTX 4500 Ada or 4-bit for consumer hardware) is frequently adopted to make large code-oriented models feasible to deploy. Collectively, CWMs establish programmatic world models as a coherent direction for code agents, coupling LLM-based program synthesis with structured, verifiable simulation for model-based reasoning.
@@ 4.3 Mathematical Agent
We note that RLVR methods such as DAPO (Yu et al., 2025e), GRPO (Ren et al., 2025), and GRESO (Zheng et al., 2025b) have consistently played a substantial role in recent enhancements of mathematical reasoning in LLMs. However, given their broader relevance across reasoning tasks, we discuss them in Section 2.7, instead of elaborating here.
@@ 4.3.1 RL for Informal Mathematical Reasoning
Informal mathematics essentially refers to reasoning and expression in natural language. Such reasoning may incorporate symbols or function names, but no finite and explicit set of logical rules defines their syntactic validity, and no formal semantics precisely determines their interpretation and meaning (Yang et al., 2024b; Asperti et al., 2025).
While informal mathematical reasoning relaxes strict rigor at the detail level, it affords greater expressive flexibility and better captures the high-level structure of arguments. This makes it particularly suited for a variety of math tasks such as mathematical word-problem solving, equation manipulation, and symbolic computation (Singh et al., 2025; Mai et al., 2025). Although general-purpose programming languages are symbolic, they lack the rigor and formal semantics of proof-assistant languages, and are therefore regarded as informal when applied to mathematical reasoning (Yang et al., 2024b), typically through tool invocation of executors such as Python with numerical or symbolic libraries.
@@
Outcome-only methods define rewards solely by final numerical or symbolic correctness (e.g., algebraic equations) during RL. Empirically, such training often leads to emergent agentic behaviors, including adaptive tool use interleaved with natural language reasoning. ARTIST (Singh et al., 2025) introduces a framework for tool-integrated agentic reasoning, interleaving tool invocations, e.g. code execution, directly within the reasoning chain. Trained with outcome-only rewards, it achieves strong performance and observes emergent agentic behaviors, including self-reflection, and context-aware CoT, which further shows that by integrating dynamic tool use with RL, agentic tool-integrated reasoning could learn optimal strategies for interacting with environments, highlighting the potential of RL to internalize tool-integrated reasoning strategies in LLMs. Similarly, ToRL (Li et al., 2025l) improves performance by exploiting the scaling of tool-integrated reasoning RL and encouraging code execution behaviour, and experiments show emergent cognitive behaviors, such as adaptive tool-use, self-correction based on tool feedback, and adaptive computational reasoning. ZeroTIR (Mai et al., 2025) investigates the scaling law of RL from outcome-based rewards for Tool-Integrated Reasoning with Python code execution settings, revealing a strong correlation between training computational effort and the spontaneous code execution frequency, the average response length, and the final task accuracy, which corroborates the empirical emergence of tool-integrated reasoning strategies. TTRL (Zuo et al., 2025) leverages majority voting to estimate rewards, enabling training on unlabeled data. Fine-tuned on these majority-vote rewards, it not only surpasses the base model’s maj@n accuracy but also achieves an empirical performance curve and upper bound that, surprisingly, closely approach those of direct RL training with labeled test answers on MATH-500, underscoring its practical value and potential. However, RENT (Prabhudesai et al., 2025) suggests that majority voting is limited in generalization, it applies only to questions with deterministic answers, and will not work on free-response outputs. To address this limitation, it extends the entropy minimization idea (Wang et al., 2021) to RL, using the token-level average negative entropy as a reward to guide learning, achieving improvements on an extensive suite of benchmarks including math problem solving, suggesting that confidence-based reward shaping can serve as a path toward continual improvement. Alternatively, Satori (Shen et al., 2025b) proposes Chain-of-Action-Thought (COAT), a variant of CoT that explicitly integrates action choices, and modularizes reasoning into 3-fold meta-actions, including continuation, reflection, and exploration of alternatives, and internalizes this behavior via RL with outcome-only rewards. In particular, 1-shot RLVR (Wang et al., 2025r) studies data efficiency of outcome-only RL with verifier signals. Surprisingly, they found that RL with only 1 example performs close to using a 1.2k-example dataset, and with 2 examples comes close to using the 7.5k MATH training dataset. They also highlight an intriguing phenomenon, named post-saturation generalization, that test accuracy continues to improve even after the training accuracy on the single example approaches 100%. In addition to correctness, hallucination remains a major challenge in informal mathematical reasoning, motivating methods that explicitly promote trustworthiness. For instance, Kirchner et al. (2024) propose a game-theoretic training algorithm that jointly optimizes for both correctness and legibility. Inspired by Prover-Verifier Games (Anil et al., 2021), the method alternates between training a small verifier that predicts solution correctness, a "helpful" prover that generates solutions accepted by the verifier, and a "sneaky" prover that aims to fool it. Empirically, this increases the helpful prover accuracy, verifier robustness and legibility (measured by human accuracy in time-constrained verification tasks). This result suggests that verifier-guided legibility optimization can enhance the interpretability and trustworthiness of LLM-generated informal reasoning. Recent rStar2-Agent (Shang et al., 2025a) is a 14B-parameter math reasoning model trained with agentic reinforcement learning using a high-throughput Python execution environment, a novel GRPO-RoC algorithm to resample on correct rollouts amid tool-noise, and a multi-stage training recipe—achieving state-of-the-art results in just 510 RL steps, achieving average pass@1 scores of 80.6% on AIME24 and 69.8% on AIME25.
@@
Process-aware methods leverage intermediate evaluators (e.g. unit tests, assertions, sub-task checks) to provide denser feedback, shaping credit assignment and improving tool-integrated reasoning (TIR). START (Li et al., 2025b) guides TIR by injecting handcrafted hint text into Long CoT traces, typically after conjunction words or before the CoT stop token, to encourage code executor calls during inference. This enables test-time scaling that improves reasoning accuracy. The collected trajectories are then used to fine-tune the model, internalizing the tool-invocation behavior. LADDER (Simonds & Yoshiyama, 2025) introduces a training-time framework where an LLM recursively generates and solves progressively simpler variants of a complex problem, using verifiable reward signals to guide a difficulty-based curriculum, and achieves substantial improvements in mathematical reasoning. An additional test-time RL step (TTRL) further enhances performance. The authors suggest that this approach of self-generated curriculum learning with verifiable feedback may generalize beyond informal mathematical tasks to any domain with reliable automatic verification. To improve performance on complex problems, SWiRL (Goldie et al., 2025) synthesizes step-wise tool use reasoning data by iteratively decomposing solutions, and then adopts a preference-based step-wise RL approach to fine-tune the base model on the multi-step trajectories. While many of these approaches exploit inference-time interventions, they often suffer from generalization limitations due to their reliance on manually designed logical structures. To overcome this, RLoT (Hao et al., 2025b) instead trains a lightweight navigator agent model with RL to adaptively enhance reasoning, showing improved generalization across diverse tasks.
While informal approaches excel at word problems and symbolic computations, they struggle to extend effectively to advanced mathematical tasks such as automated theorem proving. This limitation arises from two fundamental challenges: evaluation difficulty, which demands machine-verifiable feedback unavailable to informal methods, and scarcity of high-quality formal proof data (Yang et al., 2024b; Asperti et al., 2025).